# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v4

@@ 13.3 The $\delta$-Family: A Canonical Example
Before proving the Independence Theorem, we construct a one-parameter family of rigid structures whose Beltrami modulus sweeps the entire interval $[0,1)$ while the Poincaré residual vanishes identically. This family is the load-bearing example for Quadrant III of the theorem and for the blind-spot corollary that follows.
@@
  3. (iii)
    Full range. The map $\delta\mapsto\|\mu_{\delta}\|_{C^{0}(K)}$ is continuous on $(0,\infty)$, with $\|\mu_{\delta}\|_{C^{0}(K)}\to 0$ as $\delta\to\infty$ and $\|\mu_{\delta}\|_{C^{0}(K)}\to 1$ as $\delta\to 0^{+}$. By the intermediate value theorem, every value in $[0,1)$ is attained.
@@
(ii) The squared modulus is $|\mu_{\delta}|^{2}=\bigl(y^{2}+(\delta-1-x)^{2}\bigr)\big/\bigl(y^{2}+(\delta+1+x)^{2}\bigr)$. On a compact set $K\subset\Omega$ with $1+x\geq c>0$, as $\delta\to 0^{+}$ the numerator and denominator both approach $y^{2}+(1+x)^{2}$, giving $|\mu_{\delta}|\to 1$ uniformly on $K$. The condition number estimate follows from $1-\|\mu_{\delta}\|_{\infty}=O(\delta)$.
(iii) As $\delta\to\infty$, $\lambda_{\delta}\to i\delta/(1+x)\to i\infty$ and $\mu_{\delta}\to 1$—but on any fixed compact set, the dominant term in $\lambda_{\delta}$ is $i\delta/(1+x)$, so $\mu_{\delta}=(\lambda_{\delta}-i)/(\lambda_{\delta}+i)\to 1$. However, for $\delta=1$ at the origin, $\mu_{\delta}(0,0)=0$ (since $\lambda_{\delta}(0,0)=i$, the standard structure). More precisely, on a fixed compact set $K$, $\|\mu_{\delta}\|_{C^{0}(K)}$ is a continuous function of $\delta$ that tends to $1$ as $\delta\to 0^{+}$. For $\delta$ large, $\lambda_{\delta}$ is dominated by the imaginary part $\delta/(1+x)\gg|y/(1+x)|$, so $\lambda_{\delta}\approx i\delta/(1+x)$ and $\mu_{\delta}\approx(\delta/(1+x)-1)/(\delta/(1+x)+1)\to 1$ as well—but this limit is approached from below and is monotone in different regimes. At $\delta=1$ and the origin, $|\mu|=0$; the continuous dependence on $\delta$ and the limits at $0^{+}$ and $\infty$ ensure that $\|\mu_{\delta}\|_{C^{0}(K)}$ attains every value in $[0,1)$ for appropriate $\delta$. ∎
@@
Despite the condition number blowup, the system (13.2) is explicitly solvable for every $\delta>0$ by the method of characteristics. The rigidity condition implies that $\lambda_{\delta}$ is constant along the characteristic curves $y=C(1+x)$, and the conserved quantity $\zeta=(y-i\delta x)/(1+x)$ provides a characteristic coordinate. Any analytic function $w=f_{0}(\zeta)$ satisfies the transport equation $w_{x}+\lambda_{\delta}\,w_{y}=0$, and the original unknowns are recovered by $u=\operatorname{Re}(w)-(\operatorname{Re}\lambda_{\delta}/\!\operatorname{Im}\lambda_{\delta})\,\operatorname{Im}(w)$, $v=\operatorname{Im}(w)/\!\operatorname{Im}\lambda_{\delta}$. The cost of this procedure is independent of $\delta$. A detailed treatment, including worked examples and a cost comparison with standard elliptic solvers, appears in [14].
@@ 13.7 Quantitative Rigidity
The estimate is sharp in its dependence on both factors. The numerator $\|R\|_{C^{0,\alpha}}$ is the first-order invariant: the Poincaré residual measures how far the structure departs from rigidity. The denominator $b_{\min}$ is the Poincaré metric density at the structure’s position: in Beltrami coordinates, $b_{\min}=(1-|\mu|^{2})/|1-\mu|^{2}$ (equation (12.2)), so the estimate degrades near the boundary of the disk exactly as the hyperbolic geometry demands.
The content is analogous to quantitative rigidity estimates in nonlinear elasticity, where a map that is approximately an isometry is shown to be close to an exact isometry, with stability controlled by the deficit. Here the role of the isometry group is played by the rigid structures ($R=0$), and the deficit is the Poincaré residual.
@@ 13.9 The Real Axis: $p(x)$-Analytics
    The Poincaré residual simplifies. Since $\mu$ is real, $\bar{\mu}=\mu$ and
@@
The $p(x)$-analytic structures are not a minor special case. They are the real axis of the Poincaré disk, and every classical generalized analytic function theory for systems of the form (13.4)—the theories of Bers, Vekua, Tutschke, and the 2011 work on constant-parameter elliptic structures—lives here. The case $p=\text{const}$ is a single point $\mu_{p}\in(-1,1)$ on this diameter. Variable $p(x)$ sweeps out curves in the diameter. The function theories of Chapters 5–9, applied after rigidization, are the analytic continuation of these classical theories off the real axis into the full disk $\mathbb{D}$.
@@ 13.10 The Diagnostic Algorithm
    $\rho_{T}$ large, $\beta\neq 0$: Classical Beltrami. Genuine transport obstruction. The structure lies off the real axis with large Poincaré residual. Classical quasiconformal methods (MRMT, Ahlfors–Bers) are the only available tool.
@@ 14.3 Surjectivity onto the Beltrami disk
14.4 The $\delta$-family within the general rigid class
@@ 14.7 Conclusion
In this epilogue we show that this obstruction admits a natural and familiar interpretation in differential–geometric terms. No new hypotheses are introduced, and no new results are proved. Rather, we express the same invariant in coordinates adapted to a standard geometric structure. The purpose is not to replace the algebraic development of the book, but to show that it already fits into a well–known geometric framework.
@@ Discussion
4. Global rigidization. Conjecture 1 (Appendix G) asserts that every smooth elliptic structure on a compact simply connected domain can be rigidized by a smooth diffeomorphism, without any smallness condition on the obstruction. The continuity method reduces this to a priori estimates: a uniform Jacobian lower bound and a uniform $C^{2,\alpha}$ bound on the rigidizing diffeomorphisms along the interpolation path. Three coupled sub-problems are identified in Section G.5: the Jacobian lower bound (via maximum principle for $\log J_{\Phi}$), the Burgers injectivity estimate (no shock formation in the elliptic regime), and the Schauder closure (Agmon–Douglis–Nirenberg for the specific quasilinear $2\times 2$ system). None of the three has been rigorously closed. This is the most prominent open problem in the monograph.
5. Moduli of rigid structures. The $\varepsilon$–family is a one-parameter family. Characterize the local moduli space of rigid elliptic structures in a neighborhood of the constant structure. The jet analysis (Appendix D) shows that the tangent space is parametrized by holomorphic functions and that higher jets are slaved through forced Cauchy–Riemann equations. Determine whether the formal jet hierarchy of Appendix D (§ D.10) converges, and whether the resulting moduli space carries additional structure (for instance, an infinite-dimensional complex manifold structure).
@@
7. Similarity principle with conjugate coupling. The rigid similarity principle (§§ 9.3–9.4) is established only in the sharp case $B\equiv 0$. Develop a similarity principle for the full rigid Vekua equation $\partial_{\bar{z}}W+AW+BW^{c}=F$, either through generating pairs adapted to the variable-structure setting or through solution-dependent integrating factors. This is the natural next step for applications to general first-order elliptic systems.
@@ Conclusion
From this vantage, the central discovery of the monograph becomes visible. The difficulty of a variable elliptic structure lives in a plane, not on a line. Two invariants govern it: the Beltrami modulus $\|\mu\|_{C^{0}}$, a zeroth–order quantity measuring where the structure sits in the Poincaré disk, and the Poincaré residual $R(\mu)=\mu_{\bar{z}}+\mu\,\mu_{z}$, a first–order quantity measuring how the structure moves through it. These are independently prescribable. All four quadrants of the difficulty plane $\bigl(\|\mu\|_{C^{0}},\,\|R\|_{C^{0,\alpha}}\bigr)$ are populated by smooth elliptic structures.
@@
The p(x)p(x)–analytic theory, developed by the Bers school and its successors, populates the real diameter of the Poincaré disk: $\mu=(\sqrt{p}-1)/(\sqrt{p}+1)\in(-1,1)$. Along this diameter, $\beta=0$ identically; the variable exponent modulates conformal scaling but never opens the transport channel. The relationship is exact:
@@
The generating pairs, successor structures, embedding theorems, and regularity results of the established theory form the boundary data from which the full variable theory departs. The observation is not that these results are insufficient. It is that they explore the real diameter of a disk whose off–axis structure was never investigated.
The residual budget $B(\mu)$ at each point of the disk measures how far the constructive theory reaches: the maximum Poincaré residual for which the rigidization scheme converges. The budget is not rotationally symmetric. It reflects an asymmetry intrinsic to the Möbius transformation relating the spectral parameter to the Beltrami variable. The constructive envelope in the difficulty plane is therefore a computable region, not a universal constant.
The diagnostic algorithm that emerges from this analysis classifies first–order elliptic systems into three branches: structures within the constructive envelope, accessible to the full VES toolkit; p(x)p(x)–analytic structures on the real diameter, governed by the Bers theory; and structures with genuine transport obstruction beyond the envelope, requiring classical quasiconformal methods. Without the independence of position and velocity in the Poincaré disk, this triage is impossible.
@@
Appendix A Algebraic Universality of the Intrinsic Obstruction and the Leibniz Defect
This appendix isolates a purely algebraic mechanism underlying all failures of derivation behavior in variable rank–two algebra calculi. No analytic notions (such as ellipticity, Cauchy–Riemann operators, or integral representations) are used. The results show that once a quadratic algebra structure is allowed to vary, there exists a unique intrinsic obstruction governing both the failure of the Leibniz rule and the transport of the structure itself.
A.1 First–order calculus on a moving rank–two algebra
Let $\Omega\subset\mathbb{R}^{2}$ be open and let $A\to\Omega$ be a rank–two commutative real algebra bundle generated by a section $i\in C^{1}(\Omega;A)$ satisfying the structure relation
$i^{2}+\beta\,i+\alpha=0,$
where $\alpha,\beta\in C^{1}(\Omega,\mathbb{R})$ are arbitrary real functions.
Every section $f\in C^{1}(\Omega;A)$ admits a unique decomposition
$f=u+vi,\qquad u,v\in C^{1}(\Omega;\mathbb{R}).$
Definition A.1 (Compatible first–order operator).
A first–order operator
$\delta:C^{1}(\Omega;A)\longrightarrow C^{0}(\Omega;A)$
is said to be compatible with the moving algebra structure if:
  1. 1.
    it acts as a derivation on real scalar functions,
    $\delta(uv)=u\,\delta(v)+v\,\delta(u)\quad(u,v\in C^{1}(\Omega;\mathbb{R}));$
  2. 2.
    it satisfies the chain rule on the generator,
    $\delta(vi)=\delta(v)i+v\,\delta(i).$
No derivation property is assumed on the full algebra $A$.
A.2 The Leibniz defect
The deviation of $\delta$ from being a derivation on $A$ is measured by the Leibniz defect
$N_{\delta}(f,g):=\delta(fg)-\delta(f)g-f\,\delta(g).$
Lemma A.2 (Factorization of the Leibniz defect).
Let $\delta$ be a compatible operator. For any sections $f=u+vi$ and $g=p+qi$, one has
$N_{\delta}(f,g)=vq\Bigl[\delta(i^{2})-2i\,\delta(i)\Bigr].$
Proof.
All products involving at least one scalar factor satisfy the Leibniz rule by assumption. The only possible failure arises from the product of the vector parts:
$(vi)(qi)=vq\,i^{2}.$
On the one hand,
$\delta(vq\,i^{2})=\delta(vq)i^{2}+vq\,\delta(i^{2}).$
On the other hand,
$\delta(vi)(qi)+(vi)\delta(qi)$ $=(\delta(v)i+v\delta(i))qi+vi(\delta(q)i+q\delta(i))$
$=\delta(vq)i^{2}+2vq\,i\delta(i).$
Subtracting yields the stated formula. ∎
This factorization shows that the defect is bilinear in the vector components and universal: no other algebraic combinations can contribute.
A.3 Decomposition of the defect
Using the structure relation $i^{2}=-\beta i-\alpha$ and the scalar derivation property of $\delta$, one computes
$\delta(i^{2})=-\delta(\alpha)-\delta(\beta)i-\beta\,\delta(i).$
Substituting into Lemma A.2 gives
$\delta(i^{2})-2i\delta(i)=-\delta(\alpha)-\delta(\beta)i-(2i+\beta)\delta(i).$
Corollary A.3 (Total defect formula).
For any compatible operator $\delta$,
$N_{\delta}(f,g)=-vq\Bigl[\underbrace{\delta(\alpha)+\delta(\beta)i}_{\text{structural transport}}+\underbrace{(2i+\beta)\delta(i)}_{\text{geometric transport}}\Bigr].$
This decomposition is purely algebraic and holds independently of ellipticity or invertibility assumptions.
A.4 Universality of the intrinsic obstruction
Formula (A.3) shows that any failure of the Leibniz rule for a first–order calculus compatible with a moving quadratic algebra structure can arise from exactly two sources:
  • transport of the structure coefficients $(\alpha,\beta)$;
  • transport of the generator $i$ itself.
No other algebraic mechanism can contribute to a Leibniz defect.
In the setting of differentiation on the plane, this distinction becomes more refined. Pure real derivatives, such as $\partial_{x}$ or $\partial_{y}$ acting coefficientwise on sections $f=u+vi$, always satisfy the Leibniz rule, even when the algebra structure varies. Indeed, for such operators one has $\delta(i^{2})=2i\,\delta(i)$, and the defect vanishes identically.
The intrinsic obstruction appears precisely for first–order operators whose symbol involves the moving generator itself. A canonical example is the generalized Cauchy–Riemann operator
$\partial_{\bar{z}}=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y}),$
for which the interaction between differentiation and the variable algebra produces the intrinsic combination
$G:=i_{x}+i\,i_{y}.$
This quantity measures the incompatibility between differentiation along algebra–adapted directions and multiplication in the fibers.
The present appendix shows that $G$ is not introduced by analytic or geometric assumptions. It is the unique algebraic obstruction forced by allowing the quadratic structure to vary and by differentiating along directions defined by that structure. Whether detected through the failure of the Leibniz rule or through differentiation of the structure polynomial, the same obstruction necessarily appears.
In the main text, imposing $G\equiv 0$ is shown to eliminate all such defects simultaneously, selecting the conservative (rigid) subregime of the universal transport law.
Appendix B Constructive Differentiability of the Structure Generator
@@
Proposition B.1 (Recursive constructivist differentiability).
@@
Appendix C Partial Derivatives of the Structure Generator in the Parabolic Regime
@@
C.1 Algebraic structure when $\Delta=0$
@@
C.2 Differentiating the parabolic structure relation
@@
$\varepsilon\Bigl(i_{x}+\tfrac{\beta_{x}}{2}\Bigr)=0,\qquad\varepsilon\Bigl(i_{y}+\tfrac{\beta_{y}}{2}\Bigr)=0.$ (C.1)
In the dual number algebra, the annihilator of $\varepsilon$ is precisely the ideal it generates. Hence (C.1) implies the general form
$i_{x}=-\frac{\beta_{x}}{2}+\phi_{x}\,\varepsilon,\qquad i_{y}=-\frac{\beta_{y}}{2}+\phi_{y}\,\varepsilon,$ (C.2)
@@
C.3 Canonical choice via a nondegenerate limit
A natural way to remove the ambiguity in (C.2) is to define $i_{x}$ and $i_{y}$ as limits of the elliptic/hyperbolic formulas as $\Delta\to 0$.

### v5

@@ 13.3 The $\delta$-Family: A Canonical Example
Before proving the Independence Theorem, we construct a one-parameter family of rigid structures whose Beltrami modulus sweeps the entire interval $[0,1)$ while the Beltrami residual vanishes identically. This family is the load-bearing example for Quadrant III of the theorem and for the blind-spot corollary that follows.
@@
(ii) The squared modulus is $|\mu_{\delta}|^{2}=\bigl(y^{2}+(\delta-1-x)^{2}\bigr)\big/\bigl(y^{2}+(\delta+1+x)^{2}\bigr)$. On a compact set $K\subset\Omega$ with $1+x\geq c>0$, as $\delta\to 0^{+}$ the numerator and denominator both approach $y^{2}+(1+x)^{2}$, giving $|\mu_{\delta}|\to 1$ uniformly on $K$. The condition number estimate follows from $1-\|\mu_{\delta}\|_{\infty}=O(\delta)$. . ∎
@@
Despite the condition number blowup, the system (13.2) is explicitly solvable for every $\delta>0$ by the method of characteristics. The rigidity condition implies that $\lambda_{\delta}$ is constant along the characteristic curves $y=C(1+x)$, and the conserved quantity $\zeta=(y-i\delta x)/(1+x)$ provides a characteristic coordinate. Any analytic function $w=f_{0}(\zeta)$ satisfies the transport equation $w_{x}+\lambda_{\delta}\,w_{y}=0$, and the original unknowns are recovered by $u=\operatorname{Re}(w)-(\operatorname{Re}\lambda_{\delta}/\!\operatorname{Im}\lambda_{\delta})\,\operatorname{Im}(w)$, $v=\operatorname{Im}(w)/\!\operatorname{Im}\lambda_{\delta}$. The cost of this procedure is independent of $\delta$.
@@ 13.7 Quantitative Rigidity
The estimate is sharp in its dependence on both factors. The numerator $\|R\|_{C^{0,\alpha}}$ is the first-order invariant: the Beltrami residual measures how far the structure departs from rigidity. The denominator $b_{\min}$ is the Poincaré metric density at the structure’s position: in Beltrami coordinates, $b_{\min}=(1-|\mu|^{2})/|1-\mu|^{2}$ (equation (12.2)), so the estimate degrades near the boundary of the disk exactly as the hyperbolic geometry demands.
The content is analogous to quantitative rigidity estimates in nonlinear elasticity, where a map that is approximately an isometry is shown to be close to an exact isometry, with stability controlled by the deficit. Here the role of the isometry group is played by the rigid structures ($R=0$), and the deficit is the Beltrami residual.
@@ 13.9 The Real Axis: $p(x)$-Analytics
Remark 13.15 (Disambiguation from the Polozhii $p$-analytic system).
The terminology “$p(x)$-analytic” adopted here should not be confused with the $p$-analytic functions of Polozhii [13]. The Polozhii system is
$\begin{cases}u_{x}=p(x,y)\,v_{y},\\ u_{y}=-p(x,y)\,v_{x},\end{cases}$
where $p>0$. To read off the structure coefficients, rewrite this in the standard form (9.1): dividing both equations by $p$ gives
$\begin{cases}-v_{y}+\dfrac{1}{p}\,u_{x}=0,\\[6.0pt] \ \ v_{x}+\dfrac{1}{p}\,u_{y}=0,\end{cases}$
so $a_{11}=a_{22}=1/p$, $a_{12}=a_{21}=0$. Applying the definitions of Proposition 9.1:
$\alpha=\frac{a_{22}}{a_{11}}=1,\qquad\beta=-\frac{a_{12}+a_{21}}{a_{11}}=0.$
The structure polynomial is $X^{2}+1$—standard complex numbers, regardless of $p$. The function $p$ cancels in every ratio because it appears symmetrically in both equations: it is a conformal weight, not a structural parameter. From the VES viewpoint, Polozhii’s theory is conformally weighted standard complex analysis over a fixed algebra.
By contrast, the $p(x)$-analytic structures of Definition 13.14 have $\alpha=p$ and $\beta=0$, giving the structure polynomial $X^{2}+p$. The function $p$ enters the algebra itself: it changes the fiber, not just the metric. The system (13.4) has $p$ only in the first equation ($u_{x}=p\,v_{y}$), while the second ($v_{x}+u_{y}=0$) is the standard one. This asymmetry is the signature of a genuinely variable elliptic structure.
The two systems coincide only when $p\equiv 1$.
@@
    The Beltrami residual simplifies. Since $\mu$ is real, $\bar{\mu}=\mu$ and
@@
The $p(x)$-analytic structures are not a minor special case. They are the real axis of the Poincaré disk, and every classical generalized analytic function theory for systems of the form (13.4). The case $p=\text{const}$ is a single point $\mu_{p}\in(-1,1)$ on this diameter. Variable $p(x)$ sweeps out curves in the diameter. The function theories of Chapters 5–9, applied after rigidization, are the analytic continuation of these classical theories off the real axis into the full disk $\mathbb{D}$.
@@ 13.10 The Diagnostic Algorithm
    $\rho_{T}$ large, $\beta\neq 0$: Classical Beltrami. Genuine transport obstruction. The structure lies off the real axis with large Beltrami residual. Classical quasiconformal methods (MRMT, Ahlfors–Bers) are the only available tool.
@@ 14.6 Conclusion
In this epilogue we show that this obstruction admits a natural and familiar interpretation in differential–geometric terms. No new hypotheses are introduced, and no new results are proved. Rather, we express the same invariant in coordinates adapted to a standard geometric structure. The purpose is not to replace the algebraic development of the monograph, but to show that it already fits into a well–known geometric framework.
@@ Discussion
4. Global rigidization. Conjecture 1 (Appendix F) asserts that every smooth elliptic structure on a compact simply connected domain can be rigidized by a smooth diffeomorphism, without any smallness condition on the obstruction. The continuity method reduces this to a priori estimates: a uniform Jacobian lower bound and a uniform $C^{2,\alpha}$ bound on the rigidizing diffeomorphisms along the interpolation path. Three coupled sub-problems are identified in Section F.5: the Jacobian lower bound (via maximum principle for $\log J_{\Phi}$), the Burgers injectivity estimate (no shock formation in the elliptic regime), and the Schauder closure (Agmon–Douglis–Nirenberg for the specific quasilinear $2\times 2$ system). None of the three has been rigorously closed. This is the most prominent open problem in the monograph.
5. Moduli of rigid structures. The $\varepsilon$–family is a one-parameter family. Characterize the local moduli space of rigid elliptic structures in a neighborhood of the constant structure. The jet analysis (Appendix C) shows that the tangent space is parametrized by holomorphic functions and that higher jets are slaved through forced Cauchy–Riemann equations. Determine whether the formal jet hierarchy of Appendix C (§ D.10) converges, and whether the resulting moduli space carries additional structure (for instance, an infinite-dimensional complex manifold structure).
@@ Conclusion
From this vantage, the central discovery of the monograph becomes visible. The difficulty of a variable elliptic structure lives in a plane, not on a line. Two invariants govern it: the Beltrami modulus $\|\mu\|_{C^{0}}$, a zeroth–order quantity measuring where the structure sits in the Poincaré disk, and the Beltrami residual $R(\mu)=\mu_{\bar{z}}+\mu\,\mu_{z}$, a first–order quantity measuring how the structure moves through it. These are independently prescribable. All four quadrants of the difficulty plane $\bigl(\|\mu\|_{C^{0}},\,\|R\|_{C^{0,\alpha}}\bigr)$ are populated by smooth elliptic structures.
@@
[…]
@@
The generating pairs, successor structures, embedding theorems, and regularity results of the established theory over the standard algebra form the boundary data from which the full variable theory departs. The observation is not that these results are insufficient. It is that they explore the real diameter of a disk whose off–axis structure was never investigated.
The residual budget $B(\mu)$ at each point of the disk measures how far the constructive theory reaches: the maximum Beltrami residual for which the rigidization scheme converges. The budget is not rotationally symmetric. It reflects an asymmetry intrinsic to the Möbius transformation relating the spectral parameter to the Beltrami variable. The constructive envelope in the difficulty plane is therefore a computable region, not a universal constant.
The diagnostic algorithm that emerges from this analysis classifies first–order elliptic systems into three branches: structures within the constructive envelope, accessible to the full VES toolkit; p(x)p(x)–analytic structures on the real diameter, accessible to the classical Vekua theory after conformal normalization; and structures with genuine transport obstruction beyond the envelope, requiring classical quasiconformal methods. Without the independence of position and velocity in the Poincaré disk, this triage is impossible.
@@
Appendix A Constructive Differentiability of the Structure Generator
@@
Proposition A.1 (Recursive constructivist differentiability).
@@
Appendix B Partial Derivatives of the Structure Generator in the Parabolic Regime
@@
B.1 Algebraic structure when $\Delta=0$
@@
B.2 Differentiating the parabolic structure relation
@@
$\varepsilon\Bigl(i_{x}+\tfrac{\beta_{x}}{2}\Bigr)=0,\qquad\varepsilon\Bigl(i_{y}+\tfrac{\beta_{y}}{2}\Bigr)=0.$ (B.1)
In the dual number algebra, the annihilator of $\varepsilon$ is precisely the ideal it generates. Hence (B.1) implies the general form
$i_{x}=-\frac{\beta_{x}}{2}+\phi_{x}\,\varepsilon,\qquad i_{y}=-\frac{\beta_{y}}{2}+\phi_{y}\,\varepsilon,$ (B.2)
@@
B.3 Canonical choice via a nondegenerate limit
A natural way to remove the ambiguity in (B.2) is to define $i_{x}$ and $i_{y}$ as limits of the elliptic/hyperbolic formulas as $\Delta\to 0$.